# Adversarially Perturbed Precision Matrix Estimation

### v1

@@ Abstract
Precision matrix estimation is a fundamental topic in multivariate statistics and modern machine learning. This paper proposes an adversarially perturbed precision matrix estimation framework, motivated by recent developments in adversarial training. The proposed framework is versatile for the precision matrix problem since, by adapting to different perturbation geometries, the proposed framework can not only recover the existing distributionally robust method but also inspire a novel moment-adaptive approach to precision matrix estimation, proven capable of sparsity recovery and adversarial robustness. Notably, the proposed perturbed precision matrix framework is proven to be asymptotically equivalent to regularized precision matrix estimation, and the asymptotic normality can be established accordingly. The resulting asymptotic distribution highlights the asymptotic bias introduced by perturbation and identifies conditions under which the perturbed estimation can be unbiased in the asymptotic sense. Numerical experiments on both synthetic and real data demonstrate the desirable performance of the proposed adversarially perturbed approach in practice.
@@ 1 Introduction
The precision matrix, also known as the inverse covariance matrix, can reveal the conditional correlations between pairs of variables given all other variables (Fan et al., 2016). Estimating the precision matrix is essential in many problems, including neuroimaging (Huang et al., 2010), network analysis (Torri et al., 2018), portfolio optimization (Procacci and Aste, 2022), and genomics (Schäfer and Strimmer, 2005), among others.
Suppose $\bm{x}_{1},\ldots,\bm{x}_{n}$ are $n$ independent and identically distributed samples drawn from a $d$-dimensional multivariate normal distribution. For simplicity, we assume that the population mean is zero and the population covariance matrix $\Sigma$ is positive definite and unknown. The standard approach to obtain the precision matrix estimation, i.e., $C=\Sigma^{-1}$, is to identify the minimizer of the negative log-likelihood function:
@@
where $C\succ 0$ means that $C$ is positive definite, $\mathrm{det}C$ denotes the determinant of the matrix $C$, $\mathbb{P}_{n}$ is the empirical distribution. When the matrix dimension is small such that $d<n$, the solution to the maximum likelihood estimation problem is the inverse of the sample covariance matrix $\mathbb{E}_{\mathbb{P}_{n}}[\bm{x}\bm{x}^{\top}]$. However, if the matrix dimension is large such that $d>n$, the maximum likelihood estimation problem does not admit a solution. Numerous studies have been conducted to explore shrinkage and regularization frameworks for precision matrix estimation to achieve estimations which are invertible, well-conditioned, and enjoy desirable statistical properties such as sparsity (Yuan and Lin, 2007; Ledoit and Wolf, 2012; Nguyen et al., 2022). We will show later that the proposed adversarially perturbed framework in this paper can recover some of these shrinkage and regularized effects by selecting different types of perturbations.
@@
where $\|\cdot\|$ is the norm of the perturbation $\bm{\Delta}$, and $\delta$ is the magnitude of the perturbation. Instead of minimizing the empirical average of the likelihood function, the adversarially perturbed estimation problem (2) seeks for the minimizer of the loss under the most adversarial perturbation $\bm{\Delta}$, which is determined by both $C$ and $\bm{x}$ through the inner maximization problem in (2), subject to the constraint that the norm of the perturbation is bounded.
In this paper, we will focus on the $\ell_{p}$-norm of the perturbation. When $p=2$, it can be proven that the perturbed estimation problem (2) is equivalent to the Wasserstein shrinkage estimation proposed in Nguyen et al. (2022). The resulting optimization problem is tractable and yields a nonlinear shrinkage estimator that is well-conditioned, rotation equivalent, and preserves the eigenvalue order (Nguyen et al., 2022). When $p=\infty$, by optimizing a convex upper-bound surrogate of the objective function in the outer minimization in problem (2), the resulting estimation can be proven to promote sparsity and adversarial robustness.
We demonstrate that introducing perturbations is asymptotically equivalent to imposing regularization in precision matrix estimation, provided the perturbation magnitude is small. The corresponding first-order and second-order regularization terms are also given. Based on the regularization effect, the asymptotic distribution of the adversarially perturbed estimation is derived. The asymptotic distribution characterizes the bias induced by the inner maximization problem in our proposed framework (2). We suppose the perturbation magnitude follows the rule $\delta=\eta/n^{\gamma}$, where $\eta>0$ and $\gamma>0$. Our analysis yields the following two regimes: When $\gamma>1/2$, the asymptotic bias is zero. In this regime, the adversarial perturbation decays sufficiently fast such that it does not affect estimation accuracy. When $\gamma\leq 1/2$, the asymptotic distribution has a non-zero bias. Furthermore, specifically for $\ell_{\infty}$-norm perturbations using the upper-bound surrogate of the objective function, the derived asymptotic distribution demonstrates that sparsity recovery is achievable when $\gamma=1/2$. The regularization reformulation and asymptotic distribution provide insights into the statistical behavior of both Wasserstein shrinkage estimation and the sparse estimation recovered by our novel framework. See a summary table of the associated asymptotic behavior in Table 1, Section 5.
Numerical experiments are delivered for the case when $p=\infty$ and the surrogate objective is considered. We carry out the numerical experiments on both synthetic data and a real-world dataset. Firstly, the numerical experiments on the synthetic data validate our theoretical results that the perturbed estimation can recover the sparsity and achieve adversarial robustness. Secondly, by adopting linear discriminant analysis based on precision matrix estimation for the real dataset, our method shows more desirable practical performance.
@@ 1.1.1 Adversarial Training
As the modern artificial intelligence and machine learning models grow more powerful and widely used, their reliability becomes increasingly important. One major challenge to reliability is that machine learning models can be vulnerable to small changes in input data, which may lead to unreliable outputs (Goodfellow et al., 2015). To improve the robustness of the models, adversarial training has been introduced as a principled approach by optimizing against adversarial perturbations posed in the input data (Goodfellow et al., 2015; Madry et al., 2018).
@@
Different from the traditional empirical risk minimization approach, the resulting adversarial training problem is a min-max robust optimization problem and thus introduces a new statistical estimation paradigm. While most of the relevant existing work focus on the vector-valued parameter (Goodfellow et al., 2015; Madry et al., 2018; Javanmard and Soltanolkotabi, 2022; Ribeiro et al., 2023; Xie and Huo, 2024, 2025), our proposed framework provides novel insights on how the adversarial training motivates a new estimation framework for the matrix-valued parameter. The proposed framework serves as an alternative for practitioners when the precision matrix estimation with desirable properties such as sparsity and adversarial robustness is needed.
@@
Adversarial training is closely related to the distributionally robust optimization. It is well known that adversarial training can be viewed as a special case of the Wasserstein distributionally robust optimization (Staib and Jegelka, 2017; Gao et al., 2024). This paper, however, provides another perspective by demonstrating that, in the context of precision matrix estimation, the Wasserstein distributionally robust optimization is a special case of adversarial training.
1.1.2 Precision Matrix Estimation
@@
Precision matrix estimation has been extensively studied under different settings. Classical approaches primarily focus on enforcing sparsity via $\ell_{1}$-regularization or $\ell_{1}$-minimization (Yuan and Lin, 2007; Friedman et al., 2008; Cai et al., 2011). As applications have broadened, researchers have developed methods robust to data anomalies. For instance, Loh and Tan (2018) establishes a framework robust to data contamination, Tang et al. (2021) focuses on mitigating the impact of outliers, Avella-Medina et al. (2018) studies robustness against heavy-tailed distributions, and the method framed in Nguyen et al. (2022) addresses distributional shifts.
Our framework can recover the Wasserstein shrinkage estimation (Nguyen et al., 2022) under $\ell_{2}$-perturbation while being capable of the sparsity recovery achieved by classical $\ell_{1}$-regularized methods (Yuan and Lin, 2007; Friedman et al., 2008). In addition, by defining the model through an adversarial training lens, we achieve intrinsic adversarial robustness. Specifically, as detailed in Section 6.1, under $\ell_{\infty}$-perturbation and a specific surrogate objective, our formulation reduces to a moment-adaptive weighted $\ell_{1}$-regularized precision matrix estimation problem, thereby simultaneously achieving sparse recovery and adversarial robustness.
@@ 1.2 Notations and Definitions
$\bm{x}]_{i}$ denotes the $i$th component of vector $\bm{x}$. $\operatorname{sign}([\bm{x}]_{i})$ denotes the sign of $\bm{x}]_{i}$ (where we let $\operatorname{sign}(0)=0$), and accordingly we let $\operatorname{sign}(\bm{x})=(\operatorname{sign}([\bm{x}]_{1}),...,\operatorname{sign}([\bm{x}]_{d}))^{\top}$. We use $\|\bm{x}\|_{p}$ to denote the $\ell_{p}$-norm of the vector $\bm{z}$, i.e., $\|\bm{x}\|_{p}^{p}=\sum_{j=1}^{d}|x_{j}|^{p},1\leq p<\infty,\|\bm{x}\|_{\infty}=\max_{1\leq j\leq d}|x_{j}|.$ $\odot$ denotes the component-wise product. $\mathbb{I}(\cdot)$ denotes the indicator function. $\mathrm{vec}(X)$ denotes the vectorization of the matrix $X$.
@@ 1.3 Organization of this Paper
The remainder of this paper is organized as follows. In Section 2, we study the adversarially $\ell_{2}$-perturbed precision matrix estimation and its equivalence to the Wasserstein shrinkage estimation. In Section 3, we investigate the formulation and the properties of the adversarially $\ell_{\infty}$-perturbed precision matrix estimation. In Section 4, we introduce the regularization effect of the proposed framework. In Section 5, we study the asymptotic distribution of the proposed precision matrix estimation. Numerical experiments are conducted and analyzed in Section 6. Future work is discussed in Section 7. The proofs are relegated to the Supplementary Material whenever possible.
@@ 2 Adversarially $\ell_{2}$-perturbed Precision Matrix Estimation
where the perturbation norm in the constraint set in the inner maximization problem is chosen as $\ell_{2}$ norm.
@@ 2.1 Tractable Reformulation
which further admits the following reformulation:
$\inf_{\lambda,\{t_{i}\},C}$ $-\log\mathrm{det}C+\frac{1}{n}\sum_{i=1}^{n}\bm{x}_{i}^{\top}C\bm{x}_{i}+\lambda\delta^{2}+\frac{1}{n}\sum_{i=1}^{n}t_{i}$ (5)
@@
Problem (5) is a convex semidefinite optimization problem with a log-determinant barrier term in the objective function and linear matrix inequalities in the decision variables $\lambda,\{t_{i}\},C$. Such problems fall within the framework of semidefinite (conic) programming and can be solved efficiently by interior-point methods (Boyd and Vandenberghe, 2004; Nemirovski, 2007).
@@ 2.2 Equivalence to the Wasserstein Shrinkage Estimation
$\inf_{C\succ 0,\lambda I-C\succ 0}\left\{-\log\mathrm{det}C+\lambda\left(\rho^{2}-\frac{1}{n}\sum_{i=1}^{n}\bm{x}_{i}\bm{x}_{i}^{\top}\right)+\lambda^{2}\frac{1}{n}\sum_{i=1}^{n}\bm{x}_{i}(\lambda I-C)^{-1}\bm{x}_{i}^{\top}\right\}.$ (7)
@@
Although arriving at the same solution, our proposed perturbed framework (3) adopts a sample-wise perturbation perspective. Crucially, this formulation admits a tractable asymptotic expansion, from which the asymptotic distribution is derived, seeing Section 5. This framework elucidates the statistical effect of shrinkage and enables statistical inference.
@@ 3 Adversarially $\ell_{\infty}$-perturbed Precision Matrix Estimation
3.1 Tractable Surrogate
Notably, the inner maximization problem in problem (8) is NP-hard (Nemirovski et al., 1999). To ensure computational tractability, we replace the NP-hard inner maximization in (8) with a convex upper-bound surrogate. Based on the triangle inequality on the perturbation terms, we derive the following formulation:
$\inf_{C\succ 0}\left\{-\log\mathrm{det}C+\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{x}^{\top}C\bm{x}\right]+2\delta\sum_{j=1}^{d}\sum_{k=1}^{d}\omega_{k}|[C]_{kj}|+\delta^{2}\|C\|_{1,1}\right\}$ (9)
@@ 3.1 Tractable Surrogate
$\omega_{k}=\mathbb{E}_{\mathbb{P}_{n}}\left[\left|[\bm{x}]_{k}\right|\right].$
The formulation (9) can be interpreted as a regularized precision matrix estimation with penalties with orders $O(\delta)$ and $O(\delta^{2})$. The term $2\delta\sum_{j,k}\omega_{k}|[C]_{kj}|$ scales with $O(\delta)$ and functions as a moment-adaptive weighted $\ell_{1}$-regularizer. The weights $\omega_{k}$’s enforce a heterogeneous shrinkage scheme where entries associated with variables of larger average magnitude are penalized more heavily. Since high-magnitude features act as high-leverage points for adversarial exploitation, this weighting scheme promotes an estimation that is not only sparse but also intrinsically robust to adversarial manipulation. The quadratic term $\delta^{2}\|C\|_{1,1}$ addresses the second-order interactions of the adversarial perturbation. By penalizing the entry-wise magnitude of matrix $C$, this quadratic regularizer serves as a global stability adjustment, ensuring the surrogate remains a rigorous upper bound for the original objective in (8).
3.1.1 Properties
We will discuss the properties of the estimation obtained from (9), including the adversarial robustness stated in Proposition 3.1 and the sparsity recovery stated in Corollary 3.2.

### v2

@@ Abstract
Precision matrix estimation is a fundamental topic in multivariate statistics and modern machine learning. This paper proposes an adversarially perturbed precision matrix estimation framework, motivated by recent developments in adversarial training. The proposed framework is versatile for the precision matrix problem since, by adapting to different perturbation geometries, the proposed framework can not only recover the existing distributionally robust method but also achieve high-dimensional model selection consistency under the scale-adaptive incoherence condition, which can be viewed as a relaxation of the classic incoherence condition in the heteroscedastic settings. Additionally, the proposed perturbed precision matrix estimation framework is asymptotically equivalent to the regularized precision matrix estimation, and the asymptotic normality can be established accordingly, where the asymptotic bias introduced by perturbation is highlighted. Numerical experiments demonstrate the desirable practical performance of the proposed adversarially perturbed approach.
@@ 1 Introduction
The precision matrix, defined as the inverse of the covariance matrix, characterizes the conditional dependence structure of the data, where a zero off-diagonal entry implies that two variables are conditionally independent given all other variables (Fan et al., 2016). Estimating the precision matrix is essential in many problems, including neuroimaging (Huang et al., 2010), network analysis (Torri et al., 2018), portfolio optimization (Procacci and Aste, 2022), and genomics (Schäfer and Strimmer, 2005), among others.
Suppose $\bm{x}^{1},\ldots,\bm{x}^{n}$ are $n$ independent and identically distributed samples drawn from a $d$-dimensional multivariate normal distribution. For simplicity, we assume that the population mean is zero and the population covariance matrix $\Sigma$ is positive definite and unknown. The standard approach to obtain the precision matrix estimation, i.e., $\Sigma^{-1}$, is to identify the minimizer of the negative log-likelihood function:
@@
where $C\succ 0$ means that $C$ is positive definite, $\mathrm{det}C$ denotes the determinant of the matrix $C$, $\mathbb{P}_{n}$ is the empirical distribution. When the matrix dimension is small such that $d<n$, the solution to problem (1) is the inverse of the sample covariance matrix $\mathbb{E}_{\mathbb{P}_{n}}[\bm{x}\bm{x}^{\top}]$. However, if the matrix dimension is large such that $d>n$, the problem (1) does not admit a solution. Numerous studies have been conducted to explore shrinkage and regularization frameworks for precision matrix to achieve estimations which are invertible, well-conditioned, and enjoy desirable statistical properties such as sparsity (Yuan and Lin, 2007; Ledoit and Wolf, 2012; Nguyen et al., 2022). We will show later that the proposed adversarially perturbed framework in this paper can recover some of these shrinkage and regularization effects by selecting different types of perturbations.
@@
where $\|\cdot\|$ is the norm of the perturbation $\bm{\Delta}$, and $\delta$ is the magnitude of the perturbation. Instead of optimizing the empirical average of the likelihood function, the adversarially perturbed estimation problem (2) seeks for the minimizer of the loss under the most adversarial perturbation $\bm{\Delta}$, which is determined by both $C$ and $\bm{x}$ through the inner maximization problem in (2), subject to the constraint that the norm of the perturbation is bounded.
In this paper, we will focus on the $\ell_{p}$-norm of the perturbation.
When $p=2$, it can be proven that the perturbed estimation problem (2) is equivalent to the Wasserstein shrinkage estimation proposed in Nguyen et al. (2022). The resulting optimization problem is tractable and yields a nonlinear shrinkage estimator that is well-conditioned, rotation equivalent, and preserves the eigenvalue order (Nguyen et al., 2022).
When $p=\infty$, we optimize a convex upper-bound surrogate of the objective function in problem (2) to admit tractability. The resulting estimation is scale-adaptive and proven to achieve adversarial robustness and high-dimensional model selection consistency. The high-dimensional model selection consistency achieved by the $\ell_{1}$-regularized estimation requires the mutual incoherence condition (raskutti2008model; wainwright2009sharp; Ravikumar et al., 2011). However, our $\ell_{\infty}$-perturbed precision matrix estimation reduces the upper bound of the mutual incoherence by a factor of the ratio of the maximum to the minimum marginal standard deviation of the variables. This property facilitates consistency across a broader range of scenarios, particularly under heteroscedastic settings.
Then, we demonstrate that introducing perturbations is asymptotically equivalent to imposing regularization in precision matrix estimation, provided the perturbation magnitude is small. Based on the regularization effect, the asymptotic distribution of the adversarially perturbed estimation is derived. The asymptotic distribution characterizes the bias induced by the inner maximization problem in our proposed framework (2). We suppose the perturbation magnitude follows the rule $\delta=\eta n^{-\gamma}$, where $\eta>0$ and $\gamma>0$. Our analysis yields the following two regimes: When $\gamma>1/2$, the asymptotic bias is zero. In this regime, the adversarial perturbation decays sufficiently fast such that it does not affect estimation accuracy. When $\gamma\leq 1/2$, the asymptotic distribution has a non-zero bias. Furthermore, specifically for $\ell_{\infty}$-perturbed surrogate problem, the derived asymptotic distribution demonstrates that sparsity recovery is achievable when $\gamma=1/2$. See a summary table of the associated asymptotic behavior in Table 1, Section 5.
@@ 1.1 Related Work
As the modern artificial intelligence and machine learning models grow more powerful and widely used, their reliability becomes increasingly important. One major challenge to reliability is that machine learning models can be vulnerable to small changes in input data, which may lead to unreliable outputs (Goodfellow et al., 2015). To improve the model robustness, adversarial training has been introduced as a principled approach by optimizing against perturbations posed in the input data (Goodfellow et al., 2015; Madry et al., 2018).
@@
Different from the classic empirical risk minimization approach, the resulting adversarial training problem is a min-max robust optimization problem and thus introduces a new statistical estimation paradigm. While most of the relevant existing work focus on the vector-valued parameter in Euclidean space (Goodfellow et al., 2015; Madry et al., 2018; Javanmard and Soltanolkotabi, 2022; Ribeiro et al., 2023; Xie and Huo, 2024, 2025), our proposed framework illustrates how adversarial training inspires a novel methodology for estimating matrix-valued parameters within the positive-definite cone. To the best of our knowledge, we are the first to formulate the precision matrix estimation problem based on adversarial training.
@@
Adversarial training is closely related to distributionally robust optimization. It is well known that adversarial training can be viewed as a special case of the Wasserstein distributionally robust optimization (Staib and Jegelka, 2017; Gao et al., 2024). This paper, however, provides a reverse perspective by demonstrating that, in the context of precision matrix estimation, the Wasserstein distributionally robust optimization is a special case of adversarial training.
Precision Matrix Estimation:
Precision matrix estimation has been extensively studied under different settings. Classical approaches primarily focus on enforcing sparsity via $\ell_{1}$-regularization or $\ell_{1}$-minimization (Yuan and Lin, 2007; Friedman et al., 2008; Cai et al., 2011). The novel framework proposed in this paper is also capable of the sparsity recovery and model selection consistency. However, as mentioned in Section 1 and detailed in Section 3.2, the model consistency selection achieved by our framework is based on the scale-adaptive incoherence condition which can be weaker than that for the $\ell_{1}$-regularized estimation, especially under the heteroscedastic settings. The reason is that the adaptive incoherence in our framework has a smaller upper bound, potentially relaxing the requirements for model consistency selection when variances differ significantly.
@@
As applications have broadened, researchers have developed methods robust to data anomalies. For instance, Loh and Tan (2018) establishes a framework robust to data contamination, Tang et al. (2021) focuses on mitigating the impact of outliers, Avella-Medina et al. (2018) studies robustness against heavy-tailed distributions, and the method framed in Nguyen et al. (2022) addresses distributional shifts. From the perspective of robustness, our framework can recover the Wasserstein robust estimation (Nguyen et al., 2022) under $\ell_{2}$-perturbation. In addition, by defining the model through an adversarial training lens, we achieve intrinsic adversarial robustness. Our aforementioned scale adaptivity can also be interpreted as the robustness towards the heteroscedasticity.
@@ 1.2 Notations and Definitions
$X_{ij}$ denotes the $(i,j)$-th entry of the matrix $X$. $X_{\mathcal{I}\mathcal{J}}$ denotes the submatrix of $X$ defined by the row index set $\mathcal{I}$ and the column index set $\mathcal{J}$. $\mathrm{sign}(x)$ denotes the sign of a scalar $x$ with $\mathrm{sign}(0)=0$. Accordingly, $\mathrm{sign}(X)$ denotes the element-wise application of the sign function to matrix $X$. $\|\bm{x}\|_{p}$ denotes the $\ell_{p}$-norm of the vector $\bm{x}\in\mathbb{R}^{d}$, defined as $\|\bm{x}\|_{p}=(\sum_{j=1}^{d}|x_{j}|^{p})^{1/p}$ for $1\leq p<\infty$, and $\|\bm{x}\|_{\infty}=\max_{1\leq j\leq d}|x_{j}|$. $\|X\|_{\infty}$ denotes the induced matrix $\ell_{\infty}$-norm, defined as $\|X\|_{\infty}=\max_{1\leq i\leq d}\sum_{j=1}^{d}|X_{ij}|$. $\|X\|_{1,1}$ denotes the element-wise $\ell_{1}$-norm, defined as $\|X\|_{1,1}=\sum_{i=1}^{d}\sum_{j=1}^{d}|[X]_{ij}|$. $\|X\|_{\max}$ denotes the maximum element-wise absolute value, defined as $\|X\|_{\max}=\max_{{}_{1\leq i,j\leq d}}|X_{ij}|$. $\odot$ denotes the Hadamard product. $\otimes$ denotes the Kronecker product. $\mathbb{I}(\cdot)$ denotes the indicator function. $\mathrm{vec}(X)$ denotes the vectorization of the matrix $X$.
@@ 1.3 Organization of this Paper
The remainder of this paper is organized as follows. In Section 2, we study the adversarially $\ell_{2}$-perturbed precision matrix estimation and its equivalence to the Wasserstein shrinkage estimation. In Section 3, we investigate the formulation and the properties of the adversarially $\ell_{\infty}$-perturbed precision matrix estimation. In Section 4, we introduce the regularization effect of the proposed framework. In Section 5, we study the asymptotic distribution of the proposed precision matrix estimation. Numerical experiments are conducted and analyzed in Section 6. Future work is discussed in Section 7. The proofs are relegated to the appendix whenever possible.
@@ 2 Adversarially $\ell_{2}$-perturbed Precision Matrix Estimation
where the perturbation norm in the constraint set in the inner maximization problem is chosen as $\ell_{2}$-norm.
@@ 2.1 Tractable Reformulation
which further admits the following equivalent reformulation:
$\inf_{\lambda,\{t_{i}\},C}$ $-\log\mathrm{det}C+\frac{1}{n}\sum_{i=1}^{n}{\bm{x}^{i}}^{\top}C\bm{x}^{i}+\lambda\delta^{2}+\frac{1}{n}\sum_{i=1}^{n}t_{i}$ (5)
@@
$\begin{pmatrix}t_{i}&{\bm{x}^{i}}^{\top}C\\ C\bm{x}_{i}&\lambda I-C\end{pmatrix}\succeq 0,\quad\forall i=1,\dots,n.$
@@
Problem (5) is a convex semidefinite optimization problem with a log-determinant barrier term in the objective function and linear matrix inequalities in the decision variables $\lambda,\{t_{i}\},C$. Such problems fall within the scope of semidefinite (conic) programming and can be solved efficiently by interior-point methods (Boyd and Vandenberghe, 2004; Nemirovski, 2007).
@@ 2.2 Equivalence to the Wasserstein Shrinkage Estimation
$\inf_{C\succ 0,\lambda I-C\succ 0}\left\{-\log\mathrm{det}C+\lambda\left(\rho^{2}-\frac{1}{n}\sum_{i=1}^{n}\bm{x}^{i}{\bm{x}^{i}}^{\top}\right)+\lambda^{2}\frac{1}{n}\sum_{i=1}^{n}\bm{x}^{i}(\lambda I-C)^{-1}{\bm{x}^{i}}^{\top}\right\}.$ (7)
@@
Although arriving at the same solution, our proposed perturbed framework (3) adopts a sample-wise perturbation perspective and admits the asymptotic expansion, from which the asymptotic distribution is derived, seeing Section 4 and Section 5. In this way, our framework also helps facilitate rigorous statistical inference.
@@ 3 Adversarially $\ell_{\infty}$-perturbed Precision Matrix Estimation
The inner maximization problem in problem (8) is NP-hard (Nemirovski et al., 1999). To ensure computational tractability, we replace the NP-hard inner maximization in (8) with a convex upper-bound surrogate. Based on the triangle inequality on the perturbation terms, we derive the following formulation:
$\inf_{C\succ 0}\left\{-\log\mathrm{det}C+\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{x}^{\top}C\bm{x}\right]+2\delta\sum_{j=1}^{d}\sum_{k=1}^{d}\hat{\omega}_{k}|C_{kj}|+\delta^{2}\|C\|_{1,1}\right\},$ (9)
@@
$\hat{\omega}_{k}=\mathbb{E}_{\mathbb{P}_{n}}\left[\left|\bm{x}_{k}\right|\right].$
The formulation (9) is a regularized precision matrix. The term $2\delta\sum_{j,k=1}^{d}\hat{\omega}_{k}|C_{kj}|$ is scale-adaptive weighted $\ell_{1}$-regularizer. The weights $\omega_{k}$’s enforce a heterogeneous shrinkage scheme where entries associated with variables of larger scale are penalized more heavily. Since high-magnitude features act as high-leverage points for adversarial exploitation, this weighting scheme promotes an estimation that is not only sparse but also intrinsically robust to adversarial manipulation. The quadratic term $\delta^{2}\|C\|_{1,1}$ addresses the second-order interactions of the adversarial perturbation. By penalizing the entry-wise magnitude of matrix $C$, this quadratic regularizer serves as a global stability adjustment, ensuring the surrogate remains a rigorous upper bound for the original objective in (8). This whole regularization mechanism promotes scale-adaptive model selection and adversarial robustness, which we will be illustrated in Section 3.1 and Section 3.2.
3.1 Adversarial Robustness
We demonstrate the adversarial robustness of problem (9) in the following proposition.